# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

### v1

@@ 4.2.2 Quantifying the value of an MH connecting microtransit to Ronkonkoma station
To quantify the added value of establishing a mobility hub at Ronkonkoma for the microtransit service, we apply the same model to a network that removes the Ronkonkoma station MH (the middle hub in Fig. 3) and its MOD links. The Ronkonkoma station still remains in the dummy network. The modified network has 227 nodes and 418 links. 63 MOD service nodes and 63 MOD links remain in the MaaS platform, with only 15 census tracts losing the MOD service access via Ronkonkoma station. We obtain a solution with an optimality gap less than 0.1% in under 190 seconds. The MOD link prices are still \$2 per mile. Without the Ronkonkoma station serving as a MH, the MaaS platform usage drops from 34.19% down to only 21.4% of total travelers via the remaining two MHs. The total revenue drops to \$10,272. Therefore, establishing a MH at Ronkonkoma station with the microtransit operator is worth \$5,687 per AM period.
@@
Because the lower-level model is a coalitional choice model, the change in objective value for removing a MH captures the social surplus value of the MH.
4.2.3 Increased microtransit price rate
Because the solution obtained from the base case determines all MH based service links to be priced without subsidy, we consider the microtransit service increasing their price to \$3 per mile. This may reflect higher costs for the microtransit service such that they need to adopt a higher price rate policy.
After 195 seconds, we obtain a solution with the optimality gap less than 0.1%. This time, not all link services are priced at \$3. Instead, 25 links are priced between \$2 to \$3 per unit distance, suggesting that a subsidy is now provided for those links with less than \$3 rates. Fig. 4 illustrates the price link distribution across the network. The farthest nodes from the MHs receive lower prices, i.e. higher subsidies. This is reasonable because the longer the link, the higher the total cost. To attract more customers and generate more revenue, higher subsidies would be the most effective for longer links.
Nevertheless, the high price tolerance of the defined travelers in the area can still be identified. The total revenue is \$20195.8, a 26.5% increase from the base scenario. A total of 30% of OD flow uses the service links, which dropped from 34.19%. The extra revenue of \$4236 from the price change is only feasible with a subsidy-enabled MH design, contrary to the base scenario.
@@ 4.2.3 Increased microtransit price rate
There exists a price threshold above which a subsidy-enabled MH would outperform a traditional transit hub design.
Figure 4: Changes in link price (left) and flow (right) with increase in price cap from \$2 to \$3.
4.2.4 Flow maximization with link specific subsidy
Alternatively, the LIRR may want to consider an objective of maximizing flow instead of the base scenario. We evaluate this objective with a maximum overall subsidy of \$5000 per AM period. Eq. (14) is used to solve the problem. The algorithm terminates at the first iteration and takes 90 seconds to find a solution with the optimality gap lower that 0.1%. The final penalty term is less than $10^{-6}$, indicating a high accuracy in meeting the lower-level KKT conditions. Fig. 5 illustrates the final link prices and microtransit flows across the network. The total platform reveue is \$17,857.36\\$17,857.36. The revenue minus the \$5000 subsidy leads to a lower net revenue than the base case (by \$3101.64), but ridership increases from 34.19% (1619 passengers) up to 36.70% (1737 passengers), an increase of 118. This makes sense, as the base case shows that subsidy is not needed there, so the trade-off here is simply injecting subsidy to increase ridership at a cost to the platform.
Service areas that are further from the mobility hubs are more likely to receive subsidies, which can be observed from Fig. 5. Longer service link results in higher total service pricing, leading travelers less likely to use MOD services. Therefore, heavy subsidies are focused on further out travelers. For travelers that are closer to the hubs, they are more likely to choose hub based services even they do not receive subsidies.
Figure 5: MOD link price (left) and flow (right)
4.2.5 Maximizing flow with MH-based subsidy
We then enforce a subsidy scheme to allocate the \$5,000 subsidy uniformly for each MH. For all MOD links leading to the same mobility hub, a single subsidy value is applied. The scenario is more constrained, but reflects an alternative subsidy scheme that is dependent on the hub and more practical to control.
The proposed algorithm terminates in 45 seconds when we set the optimality target to be less than 1%, and the penalty term is also less than $10^{-6}$. However, when we set the optimality gap more aggressively to 0.1%, the algorithm fails to reach the goal before 10 iterations. Therefore, it is more computationally demanding to solve the problem using a uniform subsidy scheme. The reason is because of the increased sensitivity towards the subsidy level. When an MH based subsidy value changes, flows on all related MOD links are subject to adjust, which slows down the convergence rate. Therefore, we show the result using the one with an optimality gap of 0.63%.
Table. 5 summarizes the MH based subsidies, which are easier to compare and control than having differing levels of subsidy on every in-platform service link. Clearly the model suggests that trips to Ronkonkoma station should be subsidized 48% more than trips to St. James, likely because of the centrality of Ronkonkoma. The total MOD service link flow is 36.5% (1729 passengers). This is only a 0.5% drop compared with the link-specific subsidy allocation, but the revenue drops by 3% to \$17,288. The loss in revenue and ridership is the trade-off for having a subsidy design that is more convenient to control. Fig. 6 illustrates the differences in price and flow compared to the link-based subsidy result. The spatial distribution of subsidy differences also highlights the imbalance of subsidies with a focus towards more remote areas.
Remark 6.
MH-based subsidies are computationally more expensive than link-based subsidies but provide a subsidy design that is easier to compare and control.
Mobility hub (location in area) Subsidy level (\$/mile)
St. James (top) 0.45
Ronkonkoma (middle) 0.67
Sayville (bottom) 0.61
Table 5: Mobility hub based subsidy
Figure 6: Price difference (left) and flow difference (right) between MH-based and link-based subsidies.
@@ 5 Conclusion
In this study, we propose a MH platform design model addressing both the theoretical and computational gaps identified in recent literature for both MaaS and mobility hub based applications. While previous research has largely emphasized the role of MHs as physical transfer points or focused solely on optimal facility location, our approach extends the function of MHs to include their strategic influence on both operator and traveler decision-making, particularly through the explicit modeling of pricing structures and targeted subsidies. By leveraging the PURC framework within a bilevel optimization structure, we capture the complex interplay between public platforms and private mobility operators, while also ensuring computational scalability necessary for real-world, large-scale applications.
The proposed model captures the joint operational planning between travelers and operators needed for effective MaaS platform deployment. The bilevel framework incorporates a revenue-maximizing upper level, where the platform determines optimal pricing or subsidy allocation strategies, and a convex quadratic lower level, where travelers and operators jointly optimize flows and capacities. We transform this bilevel structure into a single-level problem using KKT conditions that guarantees the existence of a global optimum. An augmented penalty-based algorithm is proposed to improve the computational tractability and reduce runtime.
We demonstrate the model application through two sets of numerical examples. The toy example demonstrates the effectiveness of the proposed model and validates the existence of a global optimum. The model can optimally allocate link specific resources from a revenue maximization perspective. We further test the model on a large network based on LIRR stations and their service areas. A total of 244 nodes and 469 links are involved for the base network with 78 OD pairs. Several insights are obtained through the tests:
@@
    The proposed model is much more scalable compared to similar platform design models in past studies. Similar sized problems require hours of runtime ((Liu and Chow,, 2024; Yao and Zhang,, 2024)) while ours only take minutes when setting the right condition.
@@
    The use of the PURC-based lower level problem allows policymakers to quantify the social surplus value of a MH, or the value of enabling subsidy or regulating the price.
@@
    MH-based subsidies are computationally more expensive than link-based subsidies but provide a subsidy design that is easier to compare and control.
The proposed model can be applied to more complex MaaS service designs involving more operators. With more operators, subsidy schemes based on setting a rate per operator can be done. Heterogeneous travel groups can be considered (e.g. splitting each OD pair into portions by income group).

### v2

@@ 4.2.3 Quantifying the value of an MH connecting microtransit to Ronkonkoma station
To quantify the value added of the MH at Ronkonkoma, we apply the same model to a network modification where the Ronkonkoma MH (the central hub in Fig. 3) and its associated MOD links are removed. Note that the Ronkonkoma station itself remains accessible within the dummy network. In this scenario, 63 MOD service nodes and 63 MOD links remain available on the MaaS platform, with only 15 census tracts losing platform service access.
Consequently, MaaS platform usage drops significantly from 35.32% in the baseline to 22.30%, with travelers diverting to the two remaining MHs or alternative out-of-platform modes. Total platform revenue contracts to \$24,311.86, implying that the inclusion of the Ronkonkoma MH generates an incremental value of \$14,409.37 per AM peak period. Despite these volume shifts, the pricing structure remains stable. Average access fees change by less than 1%, with the first quartile shifting slightly from \$22.95 to \$23.38 and the median remaining virtually unchanged (\$25.86 vs. \$25.84).
@@
Because the lower-level model is a coalitional choice model, the change in objective value for removing a MH captures the social surplus value of the MH that travelers and operators are collectively willing to pay for, i.e. the compensating variation of the MH location alternative.
4.2.4 Lower subsidy cap provided by platform
We reduce the subsidy cap from \$5 to \$3 while keeping the original network structure to evaluate the impact of subsidy constraints on platform planning decisions. The platform profit declines to \$37,746.32, representing a 2.52% decrease compared to the baseline scenario, while the total number of travelers using the platform drops by 0.57% to 1652. Although the average access fee remains relatively stable with marginally decreasing by 1.16% to \$26.34, the allocation of subsidies shifts drastically. The total subsidy allocated to the two operators amounts to \$4,902.63 with high consolidation towards the major operator. MOD 1 sees a 1% increase in subsidy allocation with \$4,829.37, while the smaller MOD 2 faces a significant reduction of 66.81% to receiving only \$73.26. A sharp cut in subsidies caused MOD 2 ridership to fall from 149 to 42 and to only serve 15 out of the 20 covered areas, pushing the operator to the brink of market exit.
This drastic shift indicates that when subsidy flexibility is capped, the platform creates a ”flight to quality”, prioritizing the major operator (MOD 1) to secure core network capacity. The smaller operator (MOD 2), which provides a cheaper but slightly inferior service, likely relies on higher marginal incentives to keep operating. When the subsidy cap prevents the platform from bridging that gap, the platform effectively abandons the niche operator to preserve the financial viability of the primary fleet. Consequently, the restrictive cap reduces overall service diversity and accessibility, leading to a net loss in platform access and profitability. These results further reinforce the importance of model analysis of operating policies surrounding MHs, as the impacts may not be easily predicted.
@@ 4.2.4 Lower subsidy cap provided by platform
Restrictive subsidy caps can impact different operators disproportionately.
@@ 5 Conclusion
In this study, we propose a MH platform design model addressing both the theoretical and computational gaps identified in recent literature for both MaaS and mobility hub based applications. While previous research has largely emphasized the role of MHs as physical transfer points or focused solely on optimal facility location, our approach extends the function of MHs to include their strategic influence on both operator and traveler decision-making, particularly through the explicit modeling of pricing structures and targeted subsidies. By leveraging the PURC framework within a bilevel optimization structure, we capture the complex interplay between public platforms and private mobility operators while also ensuring computational scalability necessary for real-world, large-scale applications.
The proposed model integrates the joint operational planning between travelers and operators, which is essential for effective MaaS platform deployment, while also capturing the competitive dynamics of multiple operator participation. The framework utilizes a bilevel structure: a revenue-maximizing upper level, where the platform determines optimal traveler-facing pricing and operator-facing subsidy allocation strategies. A convex quadratic lower level, where travelers and operators jointly optimize flows and capacities. We transform this bilevel formulation into a single-level problem using KKT conditions, guaranteeing the existence of a global optimum. Furthermore, an augmented penalty-based algorithm is proposed to enhance computational tractability and reduce runtime.
We demonstrate the model application through two sets of numerical examples. The toy example demonstrates the effectiveness of the proposed model and validates the existence of a global optimum. The results show that the model can optimally allocate link-specific resources from a revenue-maximization perspective. We further test the model on a larger network based on LIRR stations and their service areas. A total of 380 nodes and 642 links are involved in the base network with 78 OD pairs. Several insights are obtained through the tests:
@@
    The proposed model is much more scalable compared to similar platform design models in past studies. Similar-sized problems require hours of runtime ((Liu and Chow, 2024; Yao and Zhang, 2024)), while ours takes less than 1 hour when setting the right conditions.
@@
    Platform profit exhibits a linear relationship with both MOD link operating costs and direct driving costs within the analyzed network configuration.
@@
    The platform’s profitability demonstrates greater sensitivity to changes in direct driving costs than to changes in MOD service operating costs.
  • 4.
    All three MHs demonstrate similar level of ridership shifts when subjected to variations in service operating and driving costs.
  • 5.
    Market competition is important in reducing both the overall subsidy level and platform access pricing, specifically when the operators control different levels of market power.
  • 6.
    The use of the PURC-based lower-level problem allows policymakers to quantify the social surplus value of a MH, and to determine how to control subsidies to encourage platform service participation and traveler access.
The proposed model can be applied to more complex MaaS service designs involving more operators. With more operators, subsidy schemes based on setting a rate per operator can be done. Heterogeneous travel groups can be considered (e.g. splitting each OD pair into portions by income group). The proposed model is sensitive to changes in network topology, a common feature in existing service design problems. When different spatial aggregations are used to construct the network, associated parameters, including link travel times, operating costs, and coefficients such as $\alpha_{1}$ and $\alpha_{2}$, must be adjusted and recalibrated accordingly. Therefore, proper network construction using accurate observational data is an essential prerequisite. In future studies, we aim to incorporate supply-side functions that reflect operational performance into the bi-level model. This would relax the current assumption of fixed supply-side costs, allowing the model to determine optimal operator strategies corresponding to the platform’s planning decisions.
@@ Acknowledgments
Appendix
Table A1: Simulated Demand Data
Origin (Census Tract) Destination Demand
3610300134704 NYC 37
3610300134902 NYC 75
3610300134903 NYC 50
3610300134904 NYC 72
3610300134906 NYC 44
3610300134907 NYC 56
3610300135002 NYC 76
3610300135003 NYC 57
3610300135004 NYC 65
3610300135005 NYC 31
3610300135104 NYC 60
3610300135301 NYC 52
3610300135303 NYC 87
3610300135304 NYC 97
3610300135401 NYC 46
3610300135402 NYC 35
3610300135403 NYC 83
3610300145704 NYC 47
3610300145803 NYC 41
3610300145804 NYC 90
3610300145805 NYC 64
3610300145807 NYC 78
3610300145808 NYC 52
3610300146402 NYC 83
3610300146403 NYC 83
3610300146404 NYC 72
3610300146604 NYC 53
3610300146605 NYC 49
3610300146606 NYC 58
3610300146607 NYC 62
3610300146608 NYC 40
3610300146611 NYC 67
3610300146612 NYC 42
3610300146613 NYC 109
3610300146614 NYC 67
3610300146615 NYC 64
3610300147503 NYC 72
3610300147601 NYC 87
3610300147602 NYC 58
3610300147701 NYC 47
3610300147702 NYC 28
3610300147802 NYC 63
3610300147803 NYC 47
3610300147804 NYC 57
3610300147901 NYC 58
3610300147902 NYC 44
3610300158002 NYC 76
3610300158006 NYC 72
3610300158007 NYC 29
3610300158009 NYC 33
3610300158010 NYC 51
3610300158011 NYC 29
3610300158103 NYC 72
3610300158104 NYC 48
3610300158107 NYC 77
3610300158108 NYC 87
3610300158110 NYC 97
3610300158115 NYC 55
3610300158502 NYC 58
3610300158505 NYC 53
3610300158506 NYC 78
3610300158507 NYC 67
3610300158508 NYC 62
3610300158510 NYC 31
3610300158511 NYC 65
3610300158512 NYC 56
3610300158604 NYC 78
3610300158605 NYC 51
3610300158606 NYC 62
3610300158607 NYC 67
3610300158608 NYC 74
3610300158609 NYC 17
3610300158802 NYC 75
3610300158803 NYC 46
3610300158804 NYC 64
3610300158900 NYC 68
3610300159000 NYC 59
3610300159201 NYC 72